OpenStudy (anonymous):

2 square X 20 to the power of 1/2 X 5 to the power of 3/2

OpenStudy (anonymous):

please type out your question using the equation toolbox. I can't make any sense out of your question

OpenStudy (anonymous):

2^{2} X 20_{2}^{1} X 5_{2}^{3} = ???

OpenStudy (anonymous):

it makes even less sense now. use the equation toolbox, it is a button below to the left of the post and attach file buttons.

OpenStudy (anonymous):

ok so what does \[20^{1/2}\] and \[5^{3/2}\] equal to?

OpenStudy (anonymous):

ok so what does \[20^{1/2}\] and \[5^{3/2}\] equal to?

OpenStudy (anonymous):

\[x ^{a/b} = \sqrt[b]{x^{a}}\]

OpenStudy (anonymous):

and \[\sqrt{ab}=\sqrt{a}\sqrt{b} \]

OpenStudy (radar):

\[\sqrt{2}\times \sqrt{5^{3}}\] \[\sqrt{2}\times \sqrt{5^{2}5}\] \[5\sqrt{2}\sqrt{5}=5\sqrt{10}\]